OpenStudy (anonymous):

after solving 6x^2 + 8x = 14 i was left with 3x^2+4x-7=0. after that i am unsure on how to finish it

OpenStudy (akshay_budhkar):

3x^2-3x+7x-7=0 3x(x-1)+7(x-1)=0 (x-1)(3x+7)=0 x=1 or x= -7/3

OpenStudy (akshay_budhkar):

does that help?

OpenStudy (anonymous):

where did you get the 1 from?

OpenStudy (anonymous):

did the 1 come from finding what number can be divided into both sets of numbers? 3/3=1and 7/7=1 ?

OpenStudy (akshay_budhkar):

yea exactly i took two sets of the terms in the equation 3x^2 - 3x now take 3x common 3x(x-1) same thing for the other two terms

OpenStudy (anonymous):

i understand that part now, but where did the -7/3 come from?

OpenStudy (akshay_budhkar):

(x-1)(3x+7)=0 so when we have a*b=0 then it is a=0 or b=0 so we have x-1=0 x=1 or 3x+7=0 3x=-7 x=-7/3

OpenStudy (akshay_budhkar):

is that fine?

OpenStudy (anonymous):

yeah so i just have to separate both sets of numbers and equal them to 0 and i should have my final answer?

OpenStudy (akshay_budhkar):

x= 1 or -3/7

OpenStudy (anonymous):

thank you i understand now. can you also explain how to solve (x+2)^2=1 i cant find any good examples

OpenStudy (akshay_budhkar):

ok this can be solved in another way too (x+2)^2=1 take square root on both the sides (x+2)=1 x=-1 :D

OpenStudy (anonymous):

what do you mean square root?

OpenStudy (akshay_budhkar):

\[\sqrt{x}\]

OpenStudy (anonymous):

oh! ok yeah i get it.

OpenStudy (akshay_budhkar):

\[x^2 - x +6 = 0\] i will take this example for you

OpenStudy (akshay_budhkar):

can u try it?

OpenStudy (anonymous):

is the answer x=3 or x=-2 ?

OpenStudy (katrinakaif):

At a quick glance, I Beleive that is correct

OpenStudy (anonymous):

wow im suprised i got it :D. reffering back to my second question, since x=-1 how do i find the other x? its either -3 or 4